Please show me how to graph the following parallel lines
6X +5Y = 2 and 3X + 2.5Y = -7 and how you came to this conclusion.

Answer by shree840(260)

6x+5y=2 and 3x+2.5y=-7
re write in slope intercept form y=mx+b
5y=-6x+2
y =(-6/5)x+2/5= -1.2 x+ .4
2nd eq is 2.5y=-3x-7
y= (-3/2.5)x-7/2.5
y= -1.2 x- 2.8
for parallel lines slope hase to be same so intercept will be diff.
Plot both lines